### v1

# Elliptic flow of deuterons from simulations with hybrid model

Abstract
Elliptic flow of deuterons is measured on simulated collisions events of Pb+Pb at CMS energy of 2.76 TeV per colliding nucleon pair. We use hybrid model that includes hydrodynamics for the deconfined phase and hadron transport as an afterburner. For deuterons, two production mechanisms are examined: coalescence, and direct thermal production during hadronisation and subsequent transport through the hadronic phase. Differential elliptic flow of deuterons in different centrality bins is evaluated for both implemented production models. In this approach, coalescence describes the experimental data better than direct deuteron production.
I Introduction
Production of deuterons and larger nuclear clusters from the hot matter produced in relativistic heavy-ion collisions is not fully understood, presently. The observation that their abundances are rather well described by the grand-canonical statistical model with chemical freeze-out temperature around 150 MeV Andronic:2017pug ; Andronic:2016nof is quite puzzling, as their binding energy is smaller by two orders of magnitude. This poses a question, how far one can push an interpretation with such a model in the description of other experimental observables than the multiplicities.
The other candidate model for cluster production is coalescence which binds nucleons together after their decoupling from the system, based on their proximity in phase space Butler:1963pp ; Gutbrod:1976zzr ; Sato:1981ez ; Gyulassy:1982pe ; Csernai:1986qf ; Mrowczynski:1987oid ; Lyuboshits:1988yc ; Scheibl:1998tk ; Chen:2012us ; Chen:2013oba ; Sun:2015jta ; Dong:2018cye ; Sun:2018jhg ; Sun:2018mqq ; Sombun:2018yqh ; Hillmann:2021zgj .
A recent work Braaten:2024cke offers a solution to how one can understand the success of statistical model in describing cluster multiplicity in framework of statistical physics. Nevertheless, how deuterons are produced microscopically is still an intriguing question.
Studies of deuteron production in dynamical models have long tradition Danielewicz:1991dh . More recently, SMASH hadronic transport model was used to study their production and evolution of their multiplicity during the hadronic phase Oliinychenko:2018ugs ; Staudenmaier:2021lrg , though the studies ignore the very long formation time of the deuterons Mrowczynski:2020ugu . The PHQMD model includes cluster evolution and uses a cluster finding algorithm which identifies them throughout the whole evolution of the collision Aichelin:2019tnk . The model reproduces experimental data on cluster production from nuclear collisions from few hundred MeV up to 200 GeV per colliding NN pair Glassel:2021rod ; Kireyeu:2022qmv ; Coci:2023daq ; Kireyeu:2023bye . Another kinetic model was also used for the description of light nuclei production at intermediate energies in Wang:2023gta .
In this paper we investigate with the help of a hybrid model whether differential elliptic flow can be used as a discriminator between the two candidate production mechanisms. The idea follows from the previous study Vozabova:2024jum .
The paper is structured as follows: in the next Section we explain the idea why elliptic flow could be sensitive to the production mechanism of the clusters. In Section III the setup of our simulations is explained. We used experimental data by ALICE collaboration and refer to them in Section IV. Results of our simulations are shown in Section V and the conclusions based on them are drawn in Section VI.
II Motivation
In Vozabova:2024jum , two of us have investigated the discriminating power of differential elliptic flow of clusters for the acceptance or rejection of one of the two competing production models .
Deuterons are femtoscopic probe. This is implicitly included in the formalism of coalescence and implemented since many years Mrowczynski:1987oid ; Scheibl:1998tk ; Sun:2018mqq ; Mrowczynski:1989jd ; Mrowczynski:1989jd ; Mrowczynski:1992gc ; Mrowczynski:1993cx ; Blum:2019suo ; Kachelriess:2019taq ; Kachelriess:2020amp ; Kachelriess:2023jis ; Kachelriess:2022khq In coalescence, the yield of clusters with certain momentum strongly depends on the size and density profile of the region that produces its constituent nucleons. The relevant scale to which the size is compared is the width of the cluster wave function. The yield in thermal production scales with the volume of the producing homogeneity region. The dependences on the sizes of the homogeneity regions are different in the two mechanisms.
The elliptic flow is measured as an azimuthal anisotropy of particle production. One has to recognise that particles emitted in different azimuthal directions—including clusters—are produced from different homogeneity regions, which may have different sizes. Hence, the elliptic flow of deuterons may be a sensitive probe of the production mechanism.
The results of Vozabova:2024jum have supported this conjecture. In that study, hadrons were produced from an extended blast-wave model that included second-order transverse flow and shape anisotropies, resonance production and decays, and modified freeze-out hypersurface with the freeze-out time depending on radial coordinate. The model then provided basis for two different deuteron production mechanisms. In one set of simulations, thermally produced deuterons were produced from the freeze-out hypersurface like the other hadrons. In another set of simulations, deuterons were made by coalescence. For this, nucleons were taken from the extended blast wave model, including resonance production. Then, the coalescence algorithm searched for their proximity in the phase space. The results were compared to data from Pb+Pb collisions at $\sqrt{s_{NN}}=2.76$ TeV ALICE:2015wav ; ALICE:2017nuf ; ALICE:2013mez ; ALICE:2014wao . Both transverse momentum spectra and differential elliptic flow of deuterons indicated that coalescence is the preferred production mechanism.
Here, we repeat the study with a state-of-the-art hybrid model. Such models, which combine fluid dynamics for the evolution of the deconfined phase with transport simulation for the hadronic phase, currently provide the most realistic description of ultrarelativistic heavy-ion collisions. We use a chain with TRENTo 3D for the initial conditions Ke:2016jrd , vHLLE Karpenko:2013wva for the description of the fluid in the deconfined phase, and SMASH v2.2 SMASH:2016zqf ; Karpenko:2015xea ; Schafer:2021csj as the hadronic afterburner.
Our task is to compare production of deuterons via coalescence and through direct statistical production. While the former is straightforward to implement just by using the nucleons produced in the simulation, the latter needs some explanation.
In a hybrid model, hadrons are produced statistically at the particlisation hypersurface which divides fluid-dynamical evolution and transport simulation. Then they scatter and may be annihilated as well as created in the transport phase. The observed particles are those which leave the fireball at the end of the transport phase. Their abundances and momentum distributions are finally determined there. In view of this picture, the closest scenario to the statistical production of deuterons is to generate them statistically at the hadronisation hypersurface and propagate them through the hadronic phase in transport simulation, as was done in Oliinychenko:2018ugs . This construction will be used here and compared to production by coalescence.
III The model
Our simulations use the chain TRENTo+vHLLE+Hadron Sampler+SMASH. We studied Pb+Pb collisions at $\sqrt{s_{NN}}=2.76$ TeV, for centralities up to 80%.
Initial distributions of energy density and entropy density for the fluid dynamical simulations were generated by three-dimensional version of TRENTo Ke:2016jrd . Parameters were chosen Ke:2019jbh so that the best agreement for $p_{t}$ spectra and $v_{2}(p_{t})$ is achieved for protons, pions, and kaons at all centralities, as measured by ALICE collaboration ALICE:2013mez . The values of parameters are summarised in Table 1.
Table 1: The parameters used in the simulation. TRENTo3D initial condition parameters are taken from Ke:2019jbh .
TRENTo3D
$\sqrt{s_{NN}}$ 2.76 TeV
p 0.0
k 2.0
w 0.59
n 114.0
$\mu_{0}$ 0.0
$\sigma_{0}$ 2.9
$\gamma_{0}$ 7.3
skewness 1.0
$\mathrm{J}$ 0.75
vHLLE
$\tau_{0}$ 0.6 fm/$c$
$\varepsilon_{\mathrm{crit}}$ 0.32 GeV/fm3
$\eta/s$ 0.12
$\zeta/s$ 0.4
SMASH
$\varepsilon_{\mathrm{crit}}$ 0.32 GeV/fm3
$t_{\mathrm{max}}$ 100 fm/$c$
The initial distributions were further evolved with vHLLE Karpenko:2013wva . The model assumed shear viscosity with $\eta/T=0.12$. Bulk viscosity with a maximum around critical temperature and a peak value of $\zeta/T=0.4$ was parametrised according to eq.(43) of Schenke:2020mbo , as it is also implemented in vHLLE. The fluid-dynamical simulation evolved the bulk matter down to energy density $\varepsilon_{\mathrm{crit}}=0.32$ GeV/fm3. This marked the particlisation hypersurface, at which Hadron Sampler from the vHLLE+SMASH package was used.
Each hydrodynamic evolution served as initial condition to configurations of hadrons that were further evolved with SMASH SMASH:2016zqf , version 2.2. The maximum time of the simulations was set to 100 fm/$c$.
Different centrality bins were simulated by intervals of impact parameter which were mapped to the experimentally measured percentiles of multiplicity distribution Loizides:2017ack . For each deuteron production mechanism (i.e. coalescence or direct thermal) and each centrality bin, the target number of simulated events was $1.5\times 10^{6}$. They were set up as 3000 hydrodynamic events, each oversampled by 500 SMASH afterburner runs. The simulated centrality bins were set by the finest binning from the set of the used experimental data: 0-5%, 5-10%, 10-20%, 20-30%, 30-40%, 40-50%, 50-60%, 60-70%. More peripheral events were not simulated, as clusters are truly rare there. For comparison with data organised in wider centrality bins, the corresponding originally simulated bins were merged.
For the runs where coalescence was implemented, hadrons were propagated beyond their point of last scattering or last resonance decay. Then, the same coalescence algorithm was run on (anti)protons and (anti)neutrons as in Vozabova:2024jum . First, for each proton-neutron pair their points of production or last scattering (whatever came later) were boosted into the pair rest frame. The particle produced earlier was boosted to the time when the later particle was produced. Relative momentum $\Delta p=|\boldsymbol{p}_{1}-\boldsymbol{p}_{2}|$ and distance $\Delta r=|\boldsymbol{r}_{1}-\boldsymbol{r}_{2}|$ were determined in the pair rest frame. A deuteron was created with probability 3/8 (to select the correct spin-isospin combination) if $\Delta p<\Delta p_{\mathrm{max}}$ and $\Delta r<\Delta r_{\mathrm{max}}$. The value of $\Delta r_{\mathrm{max}}=3.5$ fm was motivated by Sombun:2018yqh and $\Delta p_{\mathrm{max}}=0.310$ GeV/c was set to reproduce the observed multiplicity of deuterons. Finally, the created deuteron was boosted back to the frame in which simulation was running, to the position of the parent nucleons’ center of mass.
In runs with direct thermal production of deuterons and antideuterons, they were produced together with other hadrons during the particlisation procedure at the hadronisation hypersurface. Then, they entered the transport simulation. In SMASH, we chose the covariant collision criterion as was done in Oliinychenko:2018ugs , even though the treatment technically involves an auxiliary resonance $d^{\prime}$ with no counterpart in reality. Formally, a more consistent treatment is implemented in SMASH with the stochastic collision criterion Staudenmaier:2021lrg , but this leads to CPU times prohibitively large in view of the required statistics. For consistency, also simulations with coalescence were set with the covariant collision criterion. This is important, since a small but measurable difference in $v_{2}$ was observed between simulations with geometric and stochastic collision criterion Staudenmaier:2021lrg .
Differential elliptic flow was measured on the simulated data with the scalar product method Selyuzhenkov:2007zi , in accord with the experimental data to which we compare our results ALICE:2017nuf .
IV Experimental data
We perform our analysis on data from Pb+Pb collisions at $\sqrt{s_{NN}}=2.76$ TeV. There are most experimental data available from these collisions and we will be able to compare to the previous study Vozabova:2024jum .
In order to check the tuning of the model, we use the $p_{t}$ spectra of protons, pions, and kaons ALICE:2013mez , as well as their $v_{2}$ ALICE:2014wao . They were measured in centrality bins 0-5%, 5-10%, 10-20%, 20-30%, 30-40%, 40-50%, 50-60%.
Transverse momentum spectra of deuterons were published by ALICE collaboration in ALICE:2015wav . They used centrality bins 0-10%, 10-20%, 20-40%, 40-60%, 60-80%. Data on elliptic flow of deuterons are taken from ALICE:2017nuf . They were organised in centrality bins 0-5%, 5-10%, 10-20%, 20-30%, 30-40%, 40-50%.
V Results
Figure 1: Transverse momentum spectra of protons (left) and antiprotons (right). Different colours represent different centralities. Except for the most central data, spectra are divided by subsequent factors of $2^{n}$. Experimental data measured by the ALICE collaboration ALICE:2013mez .
Figure 2: Same as Fig. 1, but for positive (left) and negative (right) pions.
We first check that the tuning reproduces well the transverse momentum spectra and elliptic flow of hadrons. In Fig. 1 we show our results on $p_{t}$ spectra of protons and antiprotons plotted together with experimental data by ALICE collaboration ALICE:2013mez . The $p_{t}$ spectra of charged pions are shown in Fig. 2. Note that we also checked the spectra of charged kaons but omit them from showing here for brevity. The agreement between simulations and experimental data is reasonably good for $p_{t}$ below 2 GeV, where hydrodynamics for the bulk production is expected to work.
The elliptic flow of protons and pions is shown in Fig. 3.
Figure 3: Elliptic flow of identified protons and antiprotons (left) and charged pions (right). Different colors represent different centralities. Experimental data measured by the ALICE collaboration ALICE:2014wao .
The agreement between simulations and data is again reasonable. Note that it is extremely hard to reproduce the differential $v_{2}$ for all centralities and in the whole measured $p_{t}$ intervals even in much more sophisticated analyses. Even some Bayesian analyses of data with the help of some model often fit only integrated $v_{2}$ at different centralities.
These results demonstrate that the simulation is tuned appropriately and the predictions for deuteron spectra and elliptic flow can reasonably be compared to data.
Thus, we proceed with its predictions for deuterons. Their transverse momentum spectra are shown in Fig.4.
Figure 4: Transverse momentum spectra of deuterons and antideuterons. Different panels show results for different centrality bins. Dotted red line: direct thermal production of deuterons; solid blue line: deuterons built by coalescence. Experimental data by ALICE collaboration ALICE:2015wav .
Within the error intervals, both implemented mechanisms agree with the data. The direct thermal production slightly underpredicts the $p_{t}$ spectrum in the most non-central bin here: 40–60%. Nevertheless, the shape of the spectrum follows the measured one.
Our main goal, however, was to search for the difference between predictions for differential $v_{2}$. The results are shown in Fig. 5.
Figure 5: Differential elliptic flow of deuterons. Different panels show results for different centrality bins. Red dotted line: predictions from direct thermal production; blue solid line: predictions from coalescence. Bands around the lines indicate statistical uncertainties. Experimental data by the ALICE collaboration ALICE:2017nuf .
The results seem to indicate that the elliptic flow of deuterons that are thermally produced and evolved in SMASH is higher than that from coalescence. Surprisingly, this is opposite to the observation made in Vozabova:2024jum , where coalescence deuterons had higher $v_{2}$. We also note an observation which we made in our simulations, that the elliptic flow of direct thermal deuterons grows until very late times. We have noticed such a growth when comparing simulations with $t_{\mathrm{max}}=50$ fm/$c$ and $t_{\mathrm{max}}=100$ fm/$c$.
When compared to data, direct thermal production overshoots the measured $v_{2}$, while coalescence which was tuned on proton and pion production reproduces the data well.
VI Conclusions
Our simulations with hybrid model including vHLLE hydrodynamics and SMASH as hadron afterburner shows qualitatively different results for the differential elliptic flow than the previous analysis of two of us Vozabova:2024jum . Hence, we conclude that $v_{2}(p_{t})$ is qualitatively not robust enough to be the decisive probe of deuteron production mechanism.
On the other hand, our results consistently show that for both the $p_{t}$ spectra and the differential elliptic flow of deuterons, coalescence is consistent with data while the direct production of deuterons is not. Recall that the simulations were tuned on light hadrons and the description of deuterons is a prediction that follows from that tuning.
In the used model, differential elliptic flow of direct deuterons grows until very late times, when the production of light hadrons has been finished, already. Such a feature would ask for deeper study because it comes down to the question of applicability of deuteron transport through the hot hadronic has in SMASH.
The original argument which motivated this study was based on the fact that coalescence depends non-trivially in the size of the producing homogeneity region, particularly if it is comparable to the width of the cluster wave function. An interesting plan might be therefore to try the same analysis with larger clusters. There are, however, several caveats in doing this.
Firstly, scattering of larger clusters, like tritons or 4He, in a dense hadronic system is even more problematic concept by itself, as the distance between the nucleons of the cluster is comparable or larger than the typical mean free path. Such an objection could be addressed by the PHQMD model which follows the dynamics of nuclear clusters within the medium even if they are unbound Aichelin:2019tnk ; Glassel:2021rod .
Secondly, larger clusters are heavier and therefore they are suppressed in production. For example, 3He is suppressed with respect do deuterons by a factor of 300 in Pb+Pb collisions at $\sqrt{s_{NN}}=2.76$ TeV ALICE:2015wav . In turn, this means that the number of simulated events would have to be increased at least by this factor. At the moment, this is prohibitively demanding for us.
The task thus seems to be to improve the simulated statistics and look for finer probes that may be more robust and better sensitive to the production mechanism. Nevertheless, the quantitative comparison of our simulated results to data clearly prefers coalescence over direct thermal production of deuterons.
Acknowledgments
This project was supported by the Czech Science Foundation (GAČR) under Nos 22-25026S and 25-16877S. BT acknowledges support from VEGA 1/0521/22. Computational resources were provided by the e-INFRA CZ project (ID:90254), supported by the Ministry of Education, Youth and Sports of the Czech Republic. We thank Iurii Karpenko for helpful discussions.